Nursing and pharmacy students must solve dosage composition problems similar to this exercise. Set up an equation and solve by an appropriate method.
How much hydrogen peroxide in a 2% solution must be blended with hydrogen peroxide in a 20% solution to obtain 1000 liters of a 3.8% solution?

Answer by richwmiller(17219)

a+b=1000,
0.02*a+0.2*b=0.038*1000
a=1000-b
0.02*(1000-b)+0.2*b=38.0
20-0.02b+0.2*b=38.0
0.18*b=18.0
b=100.0
a=1000-b
a=900 liter at 2%
b=100 liter at 20%
check
0.02*900+0.2*100=0.038*1000
18+20=38.0
38=38.0
ok